OpenStudy (anonymous):

Use complete sentences to analyze the graph of the function f(x) = log 1/4 x. Part 1: Describe the domain, range, and general shape of this graph. (4 points) Part 2: Using complete sentences, explain how plotting specific points helps graph the function and note any critical points such as its intercepts. (4 points)

OpenStudy (anonymous):

domain is 0 to infinity range is al real numbers 

OpenStudy (anonymous):

DO you know the rest, like using complete sentences to analyze the graph?

OpenStudy (anonymous):

x intercept is 4 That specific point helps cuzz you know it will intercept there Also 1/4x due to that x cant be 0

OpenStudy (anonymous):

Alright but how do you analyze it??

OpenStudy (anonymous):

Ok but now how do you use complete sentences to analyze it?

OpenStudy (anonymous):

is the graph\[y=logx/4\] or \[y=\log_{1/4}x \]

OpenStudy (anonymous):

owleyes are you there

OpenStudy (anonymous):

Hey sorry its the first one, the 1/4 one

OpenStudy (anonymous):

Ok Thank you so much!!! I appriciate all your help:) You guys are the best!!!!!!!!

OpenStudy (anonymous):

you might say in a sentence part 1) that the graph has a domain that is more than 0 and a range is every real number,and the shape of the graph is decreasing part2)plotting specific points like (1/3;0) and asymptote x=o

OpenStudy (anonymous):

THANKSSSSS :)

OpenStudy (anonymous):

@Hero can you help me with another one of my questions? :)

OpenStudy (anonymous):

Hahaha Ok smarty pants :)

OpenStudy (anonymous):

If you say so=)

OpenStudy (anonymous):

Hahaha this is agrivating you isnt it lol

OpenStudy (anonymous):

Hahahah well then you should answer or "correct" more of my questions! lol :)

OpenStudy (anonymous):

I understand that:) Thank YOU!

OpenStudy (anonymous):

sorry about yesterday i am not in USA so the time zone is different it was around 3am so i had to sleep to start with everything the graph i was dealing with was the \[f(x)=\log_{1/4}x \] so these graph is the inverse of \[f ^{-1}(x)=(1/4)^{x}\] these is a decreasing function whose domain is \[x \in R\] and \[y>0\] so when we are dealing with inverse x and y interchange to give \[y \in R\] and \[x>0\]. However we can try out numbers for x like \[x=0,x=1/3,x=1,x=3\] then y is repectively \[y=undefined,x=1,y=0,y=-1\] proving that our graph has a range \[(\infty,-\infty)\]

OpenStudy (anonymous):

Thanks your awesome

OpenStudy (anonymous):

i still dont see which one is your question can you type it again pls

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

between \[f(x)=\log_{1/4}x \] and \[g(x)=\log_{}(x/4) \]

OpenStudy (anonymous):

First one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No thank you lol